OpenStudy (anonymous):

one week, a certain stock market index went down from 7606 to 6921.46 . what was the percent decrease, rounded to the nearest whole number?

OpenStudy (campbell_st):

what was the actual value of the drop in the stock market...?

OpenStudy (anonymous):

7606 - 6921.46

OpenStudy (anonymous):

idk

OpenStudy (campbell_st):

what ever that value you is... and you need it as percentage of the starting value \[ \frac{7606 - 6921.46}{7606} \times 100 \]

OpenStudy (anonymous):

684.54

OpenStudy (campbell_st):

nope.... when you do the calculation, put the numerator in brackets.

OpenStudy (anonymous):

idk what your sayin im only in 7th grd

OpenStudy (campbell_st):

do you have a calculator

OpenStudy (anonymous):

yea

OpenStudy (campbell_st):

ok... this is what you enter \[(7606 - 6921.46)\div7606 \times 100 = \] tell me what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

um i got 9

OpenStudy (campbell_st):

the steps to answering the question are 1. find the difference 7606 - 6921.46 = 684.54 2. create a fraction: difference/starting value \[\frac{684.54}{7606}\] 3. to change a fraction to a percentage, multiply by 100. \[\frac{684.54}{7606} \times 100\]

OpenStudy (campbell_st):

thats what I go... well done...

OpenStudy (anonymous):

thx